Hello! Can you please solve this problem because I'm really having a hard time understanding this problem, pls.....I really need this.
This is all about distance/rate problem.
1. Two different routes between two towns differ by 24 kilometers. One motorist traveled the longer route at 65km/hr and another traveled the shorter route at 52 km/hr. If they made the trip between two towns in exactly the same time, what is the length of each route?
P.S.
I've tried answering it and I came up with route A = 120.25 ; route B 96.2
but I don't know if this is correct.
>>Thank You<<

Answer by jorel1380(3719)

65n=52n+24
13n=24
n=24/13
65x24/13=120
52x24/13=96
120-96=24
They both took 24/13 hours, or 1.8462 hours to reach their destination. You were close, but the actual routes were 120 and 96 kilometers, respectively..